# Local Level Dynamic Random Partition Models for Changepoint Detection

### v1

@@ 4.1 Simulations with independent data
We implemented the Gibbs sampling scheme that was detailed in Section 3. We ran the algorithm for 10,000 iterations, with the first half discarded as burn-in. Subsequently, we calculated the optimal partition for each time point based on the estimated posterior similarity matrix and the estimated posterior probability $\eta_{t}$, for $t=1,\ldots,T$. To determine whether a specific time point $t$ should be considered a changepoint, we employed the Bayesian False Discovery Rate method (Newton et al., 2004; Müller et al., 2006). As explained in Section 3, we implemented a penalized version of FDR with a control level set at (0.01/3).
Table 1 presents the performance of the five models across varying data dimensions. Remarkably, the LLDPM consistently outperforms all other models across all metrics. However, as the sample size increases, all models show a decrease in performance due to the higher likelihood of subjects switching clusters. The results underscore the robust performance of GMDDP in metrics related to changepoint detection. At the same time, when evaluating the accuracy of recovering the true latent cluster structure, depicted in Figure 3, GMDDP struggles and does not accurately identify clusters. Therefore, GMDDP may not be optimal when simultaneous changepoint detection and cluster analysis are required. Also the DRPM shows robust performance in changepoint detection, with an average specificity comparable to our model in each scenario. However, its ability to recover the correct cluster structure diminishes as the number of subjects increases. In contrast, LLDPM demonstrates excellent performance in both changepoint detection and recovering the latent cluster structure.
@@
Figure 3: Section 4.1. Boxplots of the average ARI, to evaluate the clustering performance when analyzing independent data, for our model and four alternatives, with $n=\{20,50,100\}$. Results are based on 50 replicated datasets and averaged over 100 time points. For each dataset and each method, at each time point, the true partition is compared with the partition estimated by minimizing the lower bound of the posterior expected Variation of Information.
@@ 4.2 Simulations with autoregressive data
$\lambda$ Measure LLDPM DRPM LDDP WDDP GMDDP 0.25 specificity 0.8683 (0.151) 0.8728 (0.152) 0.9971 (0.014) 0.8121 (0.066) 0.7700 (0.125) accuracy 0.9147 (0.059) 0.5156 (0.044) 0.9987 (0.006) 0.5029 (0.032) 0.6173 (0.068) recall 0.9175 (0.073) 0.2031 (0.097) 1.000 (0.000) 0.2323 (0.063) 0.4838 (0.065) precision 0.9252 (0.065) 0.7473 (0.211) 0.9976 (0.012) 0.6043 (0.132) 0.7219 (0.115) F1 0.9195 (0.055) 0.2972 (0.095) 0.9988 (0.006) 0.3280 (0.078) 0.5743 (0.064) AUC 0.9145 (0.059) 0.5284 (0.040) 0.9986 (0.007) 0.5204 (0.031) 0.6308 (0.063) 0.5 specificity 0.9082 (0.177) 0.9059 (0.178) 0.9943 (0.020) 0.8111 (0.061) 0.5514 (0.198) accuracy 0.8913 (0.066) 0.5276 (0.039) 0.9973 (0.009) 0.5015 (0.036) 0.5800 (0.100) recall 0.9038 (0.082) 0.1966 (0.134) 1.000 (0.000) 0.2306 (0.070) 0.6050 (0.104) precision 0.8992 (0.074) 0.8368 (0.208) 0.9953 (0.016) 0.5772 (0.085) 0.6219 (0.111) F1 0.8984 (0.061) 0.2886 (0.109) 0.9976 (0.008) 0.3249 (0.088) 0.6059 (0.082) AUC 0.8904 (0.066) 0.5421 (0.044) 0.9971 (0.010) 0.5188 (0.034) 0.5866 (0.097) 0.75 specificity 0.9226 (0.081) 0.9286 (0.066) 0.9671 (0.065) 0.7726 (0.071) 0.3514 (0.299) accuracy 0.8707 (0.060) 0.5352 (0.036) 0.9847 (0.030) 0.4993 (0.061) 0.6533 (0.156) recall 0.8838 (0.082) 0.1910 (0.077) 1.000 (0.000) 0.2602 (0.122) 0.9175 (0.092) precision 0.8862 (0.089) 0.7940 (0.168) 0.9747 (0.049) 0.5618 (0.089) 0.6383 (0.139) F1 0.8797 (0.062) 0.2972 (0.096) 0.9866 (0.026) 0.3474 (0.102) 0.7458 (0.107) AUC 0.8697 (0.062) 0.5430 (0.040) 0.9836 (0.032) 0.5161 (0.058) 0.6345 (0.164) 0.9 specificity 0.9127 (0.095) 0.9143 (0.096) 0.4586 (0.296) 0.7500 (0.094) 0.2811 (0.269) accuracy 0.8533 (0.056) 0.5181 (0.054) 0.7407 (0.141) 0.4680 (0.052) 0.6290 (0.133) recall 0.8600 (0.078) 0.1714 (0.091) 0.9875 (0.028) 0.2213 (0.066) 0.9334 (0.157) precision 0.8701 (0.072) 0.7380 (0.223) 0.6977 (0.130) 0.5090 (0.131) 0.6020 (0.149) F1 0.8618 (0.053) 0.2648 (0.123) 0.8112 (0.089) 0.3041 (0.081) 0.7420 (0.083) AUC 0.8529 (0.057) 0.5300 (0.050) 0.7230 (0.150) 0.4856 (0.053) 0.5987 (0.136)
@@
Table 2: Section 4.2. Performance measures for changepoint detection with AR(1) data, for the LLDPM and four competing models. The values correspond to the average (standard errors) over 50 simulations.
@@
In a second simulation study, we aim to assess the performance of our LLDPM in scenarios where the data are generated from a process characterized by an autoregressive structure of order one, or AR(1). Specifically, we consider datasets with $T=30$ time points and $n=20$ units, generated from the model
$Y_{i,t}=\lambda Y_{i,t-1}+\tilde{\beta}_{i,t}+\tilde{\varepsilon}_{i},$
@@
where the error terms $\tilde{\varepsilon}_{i}$ are assumed to be independent and identically distributed from a Normal with mean zero and variance one. At each time point $t=1,\ldots,T$, units are partitioned based on the distinct values taken by $\tilde{\beta}_{i,t}$. The number of distinct values of $\tilde{\beta}_{i,t}$, and hence the number of clusters, ranges between one and three. Specifically, it is equal to three if $t$ is divisible by five, it is equal to two if $t$ is divisible by nine, and it is equal to one otherwise. The individual allocation of the units to clusters changes every time the number of clusters changes. An example of such data is shown in Figure C.9 in the Supplementary Material. When fitting our model we set $\theta$ according to the resulting expected prior number of cluster, which for the LLDPM with specification $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$ is given by $\mathbb{E}[|\pi_{t}|]=\sum_{i=1}^{n}\theta/(\theta+i-1)$ (Pitman, 2002), for every $t=1,\ldots,T$. We thus set $\theta=0.32$, corresponding to $\mathbb{E}[|\pi_{t}|]=2$. For the other parameters, we use the same specifications as in the first simulation study (see Section 4.1). As before, we run the MCMC for 10,000 iterations, discarding the first half.
The results of our analysis, based on 50 replicates, are presented in Table 2. The LLDPM consistently excels in identifying changepoints across all metrics considered. In this second scenario, the GMDDP shows reduced effectiveness compared to the previous simulation study. The LDDP demonstrates strong ability in identifying changepoints; however, similar to the GMDDP in the previous study, it is less reliable in accurately identifying the cluster structure. Figure 4 illustrates that our LLDPM outperforms the other methods in recovering the correct latent cluster allocations, with ARI indices consistently above 0.7 for data generated with different autoregressive coefficients.
Considering both scenarios together, our approach appears to provide reliable inference both for changepoint detection and cluster allocation with respect to the competing approaches considered here.
@@
Figure 4: Section 4.2. Boxplots of the average ARI to evaluate the clustering performance when analyzing AR(1) data, for our model and four alternatives, and different values of the autoregressive coefficient $\lambda=\{0.25,0.5,0.75,0.9\}$. Results are based on 50 replicated datasets and averaged over 30 time points. For each dataset and each method, at each time point, the true partition is compared with the partition estimated by minimizing the lower bound of the posterior expected Variation of Information.
5 Application to Gesture Phase Segmentation
@@
We present an analysis of video-recorded data for human gesture segmentation. The goal is to segment videos into distinct phases exhibiting different motion patterns, e.g., to identify time lapses of the video that need to be removed from a clip (Parvathy et al., 2021). More specifically, we employ the Gesture Phase Segmentation dataset, originally described by Madeo et al. (2013), which is publicly accessible for download at the following URL: https://archive.ics.uci.edu/ml/datasets/gesture+phase+segmentation. While Madeo et al. (2013) analyzed this dataset using Support Vector Machines (SVM) to segment, and classify gesture data streams, they did not consider partitioning gesture phases into their components and investigating the temporal dependence between partitions. Our goal, in contrast, is to detect the latent cluster structure at each frame while capturing changepoints, using our proposed LLDPM random partition model with time dependence.
@@
The dataset contains sensor data recordings of users recounting comic book stories facing an Xbox Microsoft KinectTM sensor. The dataset provides scalar velocity and acceleration values over four sensors, placed on the left hand, right hand, left wrist, and right wrist, leading to $n=8$ sensor measurements at regular time intervals (frames). These values were obtained by normalizing hand and wrist positions relative to the head and spine position using a fixed displacement offset of 3 to measure velocity. Our analysis is based on a processed version of the data, which we prepare following the approach outlined by Hadj-Amar et al. (2024). The preprocessing involves four stages:
@@
Figure 5: Section 5, Human Gesture data. Scalar velocity of the left and right hand and the wrists after preprocessing ($T=349$).
@@
Figure 6: Section 5, Human Gesture data. Changepoints (dashed line) identified with the LLDPM.
@@
The estimated changepoints are displayed in Figure 6 and appear distributed throughout the time series, possibly indicating that the person wearing the sensors is gesticulating while recalling the comic story. When compared with the segmentation in Madeo et al. (2013), the changepoints identified by the LLDPM primarily occur during the activity phase, which correspond to the preparation, retraction, and stroke phases in the video. Additionally, most of these changepoints are observed during transitions between different phases, with a notable increase in frequency during the reading phases. These changepoints likely reflect movements related to the subject’s gesticulation and body language while narrating, as well as shifts in hand and wrist positions during the assigned activity. Figure 7 illustrates the clustering of the eight sensors within a specific time window, chosen for representation purposes. The estimated clustering is obtained by minimizing the lower bound of the posterior expected Variation of Information. The numbers on the plot correspond to the sensor measurements. Notably, a distinct pattern allocation emerges, where all the sensors are grouped into a single primary cluster. Additionally, a secondary pattern takes shape, forming two clusters: one containing all the sensors from the left arm (odd numbers) and the other with sensors from the right arm (even numbers) grouped together.
@@
For comparison, we analyzed the same data with the alternative models considered in this work. The identified changepoints are displayed in Figures, D.11, D.12, D.13 and D.14 in the Supplementary Material. It is notable that partition-based models appear to identify fewer changepoints. For instance, the DRPM detected only 15 changepoints (4.3% of the total time points), primarily at the beginning of the time series, and the LLDPM identified 23 changepoints (6.6%). In contrast, the other alternative models, namely the LDDP, WDDP, and GMDDP, identified approximately 14% of the changepoints, with 48, 50, and 51 changepoints, respectively, out of 349 time points. These observations instil confidence in the performance of the LLDPM, which thus might a reliable modeling option, particularly in real-world applications where the ground truth is unknown.
@@
Figure 7: Section 5, Human Gesture Data. Time-specific clusters obtained by minimizing the lower bound of the posterior expected Variation of Information, for the time window from time 335 to time 340. Numbers refer to the measurements of scalar velocity (1-4) and acceleration (5-8) in sensor units placed on the hands and wrists of a subject, whereas colors indicate clusters.

### v2

@@ 4.1 Simulations with independent data
We implemented the Gibbs sampling scheme described in Section 3, running the algorithm for 10,000 iterations and discarding the first 5,000 as burn-in. We randomly selected a subset of replications and visually inspected the trace plots of the time-specific number of clusters and the partition entropy. These diagnostics revealed no signs of convergence issues. Examples of these trace plots are provided in Section C.2 of the Supplementary Material. Subsequently, we calculated the optimal partition for each time point based on the estimated posterior similarity matrix. To determine whether a specific time point $t$ should be considered a changepoint, we employed the Bayesian False Discovery Rate method (Newton et al., 2004; Müller et al., 2006). As explained in Section 3, we implemented a penalized version of FDR with a control level set at 0.01/3.
Table 1 summarizes the performance of the seven models considered, including our LLDPM0 and LLDPM0.25, and models 1–5 introduced at the beginning of Section 4, across varying data dimensions $n$. Notably, both versions of our model exhibit strong performance, consistently outperforming LDDP, WDDP, and PPM. All changepoint detection metrics highlight the robustness of LLDPM, displaying strong and stable results regardless of the underlying base process specification, whether based on a CRP or a 2-CRP with $\sigma=0.25$. Further remarks are needed for the remaining two competitors, GMDDP and DRPM. While GMDDP performs well in detecting changepoints, as reflected in the evaluation metrics, it struggles to accurately recover the true time-specific latent cluster structures, as illustrated in Figure 3. Consequently, GMDDP may not be well-suited for applications where both changepoint detection and cluster analysis are of interest. Also DRPM shows strong results in changepoint detection, achieving specificity levels comparable to those of LLDPM across all settings. However, it suffers from lower recall, indicating a tendency to miss many true changepoints. In contrast, LLDPM excels in both tasks, consistently delivering high changepoint detection accuracy and near-perfect ARI scores for latent cluster recovery, see Figure 3. Finally, we observe that as the sample size increases, performance tends to decline across all models, due to the greater likelihood of subjects switching clusters.
To evaluate the performance of our model for sample sizes larger than those considered in Table 1, we used both versions of the LLDPM to analyze 50 replicated datasets with $n=500$ observations over $T=100$ time points. Due to memory limitations and prohibitively long runtimes, the competing models could not be fitted. In contrast, for our models, we adopted a blockwise Gibbs sampling strategy to manage the computational load, splitting the sampling into smaller blocks of iterations to prevent memory and performance issues. Both LLDPM0 and LLDPM0.25 successfully identified the changepoints, achieving perfect recall. However, their performance in correctly identifying periods with no-changepoint declined, leading to an increase in false-positives. Specifically, LLDPM0 achieved over 70% specificity and accuracy, with an AUC close to 80%. However, its F1 score dropped to around 40% due to lower precision. LLDPM0.25 showed a more polarized behavior with specificity and accuracy around 40%, an AUC of 70%, and an F1 score of about 25%. Despite these differences in changepoint detection metrics, both models achieved an average ARI of approximately $99.95\%$, indicating a near-perfect recovery of the time-specific latent cluster structures.
Figure 3: Section 4.1. Boxplots of the average ARI, to evaluate the clustering performance when analyzing independent data, with $n=\{20,50,100\}$, for two versions of our LLDPM model and four alternatives. The PPM is not included in the comparison as the model does not account for time-specific partitions of the units
@@
. Results are based on 50 replicated datasets and averaged over 100 time points. For each dataset and each method, at each time point, the true partition is compared with the partition estimated by minimizing the lower bound of the posterior expected VI.
@@ 4.2 Simulations with autoregressive data
$\lambda$ Measure LLDPM0 LLDPM0.25 DRPM LDDP WDDP GMDDP PPM AR1-DP 0.25 specificity 0.91 (0.08) 0.95 (0.06) 0.98 (0.04) $\mathbf{1.00}$ (0.00) 0.82 (0.05) 0.87 (0.10) 0.83 (0.02) 0.64 (0.15) accuracy 0.91 (0.06) 0.94 (0.04) 0.52 (0.04) $\mathbf{1.00}$ (0.02) 0.52 (0.03) 0.65 (0.07) 0.49 (0.00) 0.83 (0.07) recall 0.92 (0.07) 0.94 (0.06) 0.13 (0.07) 1.00 (0.04) 0.25 (0.05) 0.45 (0.07) 0.21 (0.02) $\mathbf{1.00}$ (0.00) precision 0.93 (0.07) 0.96 (0.05) 0.91 (0.16) $\mathbf{1.00}$ (0.00) 0.62 (0.09) 0.81 (0.12) 0.61 (0.01) 0.77 (0.08) F1 0.92 (0.06) 0.95 (0.04) 0.21 (0.11) $\mathbf{1.00}$ (0.02) 0.35 (0.06) 0.58 (0.08) 0.31 (0.02) 0.87 (0.05) AUC 0.91 (0.06) 0.94 (0.04) 0.54 (0.04) $\mathbf{1.00}$ (0.02) 0.53 (0.03) 0.66 (0.07) 0.52 (0.00) 0.82 (0.08) 0.5 specificity 0.88 (0.10) 0.94 (0.07) 0.96 (0.18) $\mathbf{0.99}$ (0.02) 0.80 (0.06) 0.67 (0.16) 0.82 (0.02) 0.55 (0.14) accuracy 0.89 (0.07) 0.93 (0.05) 0.51 (0.04) $\mathbf{1.00}$ (0.01) 0.51 (0.04) 0.58 (0.08) 0.49 (0.00) 0.79 (0.07) recall 0.90 (0.08) 0.93 (0.07) 0.13 (0.13) $\mathbf{1.00}$ (0.00) 0.25 (0.06) 0.51 (0.07) 0.22 (0.03) $\mathbf{1.00}$ (0.00) precision 0.90 (0.07) 0.95 (0.06) 0.84 (0.21) $\mathbf{0.99}$ (0.02) 0.59 (0.08) 0.65 (0.12) 0.60 (0.01) 0.72 (0.07) F1 0.90 (0.06) 0.94 (0.05) 0.21 (0.11) $\mathbf{1.00}$ (0.01) 0.35 (0.07) 0.57 (0.07) 0.32 (0.02) 0.84 (0.04) AUC 0.89 (0.07) 0.93 (0.05) 0.54 (0.04) $\mathbf{1.00}$ (0.01) 0.52 (0.04) 0.60 (0.07) 0.52 (0.00) 0.78 (0.07) 0.75 specificity 0.86 (0.12) 0.92 (0.08) 0.94 (0.07) $\mathbf{0.96}$ (0.08) 0.77 (0.07) 0.45 (0.25) 0.83 (0.02) 0.34 (0.14) accuracy 0.87 (0.06) 0.92 (0.06) 0.52 (0.04) $\mathbf{0.98}$ (0.04) 0.49 (0.04) 0.66 (0.16) 0.49 (0.00) 0.69 (0.07) recall 0.88 (0.08) 0.91 (0.08) 0.16 (0.08) $\mathbf{1.00}$ (0.00) 0.24 (0.06) 0.84 (0.14) 0.22 (0.02) 1.00 (0.01) precision 0.89 (0.09) 0.93 (0.07) 0.79 (0.17) $\mathbf{0.97}$ (0.06) 0.55 (0.09) 0.65 (0.14) 0.61 (0.01) 0.64 (0.05) F1 0.88 (0.06) 0.92 (0.06) 0.25 (0.10) $\mathbf{0.98}$ (0.03) 0.33 (0.07) 0.73 (0.12) 0.32 (0.02) 0.78 (0.04) AUC 0.87 (0.06) 0.92 (0.06) 0.53 (0.04) $\mathbf{0.98}$ (0.04) 0.51 (0.04) 0.65 (0.16) 0.52 (0.00) 0.67 (0.07) 0.9 specificity 0.85 (0.10) 0.90 (0.10) $\mathbf{0.95}$ (0.10) 0.49 (0.27) 0.77 (0.09) 0.37 (0.31) 0.81 (0.03) 0.22 (0.10) accuracy 0.87 (0.06) $\mathbf{0.90}$ (0.06) 0.51 (0.04) 0.75 (0.13) 0.49 (0.04) 0.65 (0.16) 0.49 (0.01) 0.62 (0.05) recall 0.88 (0.06) 0.90 (0.07) 0.13 (0.08) $\mathbf{0.98}$ (0.04) 0.24 (0.07) 0.91 (0.19) 0.23 (0.03) 0.97 (0.04) precision 0.89 (0.08) $\mathbf{0.91}$ (0.07) 0.79 (0.17) 0.70 (0.12) 0.55 (0.12) 0.63 (0.18) 0.60 (0.01) 0.59 (0.03) F1 0.88 (0.07) $\mathbf{0.90}$ (0.05) 0.22 (0.10) 0.81 (0.08) 0.33 (0.08) 0.75 (0.12) 0.33 (0.03) 0.73 (0.03) AUC 0.86 (0.06) $\mathbf{0.90}$ (0.06) 0.52 (0.05) 0.73 (0.14) 0.50 (0.04) 0.65 (0.15) 0.52 (0.00) 0.60 (0.05)
@@
Table 2: Section 4.2. Performance measures for changepoint detection with AR(1) data for two versions of our LLDPM model and six alternatives. Values correspond to the average (standard error) over 50 replicates.
@@
In a second simulation study, we aim to assess the performance of the two versions of our LLDPM in scenarios where the data are generated from a process characterized by an autoregressive structure of order one, or AR(1). Specifically, we consider datasets with $T=30$ time points and $n=20$ units, generated from the model
$Y_{i,t}=\lambda Y_{i,t-1}+\tilde{\beta}_{i,t}+\tilde{\varepsilon}_{i,t},$ (19)
@@
where the error terms $\tilde{\varepsilon}_{i,t}$ are assumed to be independent and identically distributed from a Normal with mean zero and variance one. At each time point $t=1,\ldots,T$, units are partitioned based on the distinct values taken by $\tilde{\beta}_{i,t}$. The number of distinct values, and thus the number of clusters, varies between one and two over time. Specifically, there are two clusters when $t$ is divisible by five or nine, and one cluster otherwise. When $t$ is divisible by five, the two clusters are of equal size; when $t$ is divisible by nine, one cluster is more than twice the size of the other. This setup gives rise to three distinct configurations that alternate across time points, with the allocation of units to clusters updated whenever the configuration changes. The observations are thus simulated from model (19), with the values of $\tilde{\beta}_{i,t}$ at each time point assigned to arbitrarily fixed, cluster-specific values. An example of such data is shown in Figure C.12 in the Supplementary Material. When fitting the two versions of our model, unlike in Section 4.1, we set $\theta$ based on the expected prior number of clusters, $\mathbb{E}[|\pi_{t}|]$, for each time point $t=1,\ldots,T$. For the LLDPM with specification $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$, this is given by $\mathbb{E}[|\pi_{t}|]=\sum_{i=1}^{n}\theta/\theta+i-1$, while for the LLDPM with $p^{*}(\cdot)=p_{\sigma}(\cdot)$ and $\sigma\in(0,1)$, by $\mathbb{E}[|\pi_{t}|]=\left[(\theta+\sigma)_{n}/(\theta+1)_{n-1}-\theta\right]/\sigma$ (Pitman, 2006). By setting $\mathbb{E}[|\pi_{t}|]=2$ for all $t=1,\ldots,T$, we obtain $\theta=0.32$ for LLDPM0 and $\theta=-0.07$ for LLDPM0.25. Moreover, we let the autoregressive parameter $\lambda$ vary in $\{0.25,0.5,0.75,0.9\}$ in order to see if increasing the dependence across times affect the performances of the model. For the other parameters, we use the same specifications as in the first simulation study (see Section 4.1). As before, we run the MCMC for 10,000 iterations, discarding the first half as burn-in.
The results of our analysis, based on 50 replicates, are presented in Table 2. Both versions of the LLDPM consistently outperform most other models in detecting changepoints across all considered metrics. In this second study, we observe that GMDDP exhibits reduced effectiveness compared to the previous simulation. Among the alternatives, LDDP stands out for its strong ability to identify changepoints, although its performance declines in terms of specificity, precision, and accuracy as $\lambda$ increases. In contrast, both versions of the proposed LLDPM demonstrate robustness, maintaining high performance across all metrics and scenarios, regardless of the degree of temporal dependence in the simulated data. This consistency highlights the reliability of the LLDPM across a wide range of settings. Furthermore, the LDDP falls short in accurately recovering the time-specific latent cluster structure. As shown in Figure 4, both versions of the LLDPM consistently achieve ARI indices above 0.7 across datasets generated with varying autoregressive coefficients, clearly outperforming the other methods. Overall, when considering both changepoint detection and cluster structure recovery, our model provides reliable and robust inference across a wide range of scenarios.
Figure 4: Section 4.2. Boxplots of the average ARI are presented to evaluate clustering performance when analyzing AR(1) data, comparing two versions of our model with five alternatives. The PPM is excluded from the comparison as it does not account for time-specific partitions of the units.
@@
The analysis considers different values of the autoregressive coefficient $\lambda=\{0.25,0.5,0.75,0.9\}$. Results are based on 50 replicated datasets and averaged over 30 time points. For each dataset and each method, at each time point, the true partition is compared with the partition estimated by minimizing the lower bound of the posterior expected VI.
5 Applications
We apply our LLDPM model, using the specification $p^{*}(\cdot)=p_{\text{CRP}}(\cdot)$, to two real-world datasets: first, the Gesture Phase Segmentation data (Madeo et al., 2013), and second, data from the Collaborative Perinatal Project (CPP, Longnecker et al., 2001). Gesture Phase Segmentation focuses on identifying and classifying different phases of hand movements, which is an important task in applications such as sign language recognition, human-computer interaction, and motion analysis. In contrast, CPP is a large-scale epidemiological study that investigates prenatal and perinatal factors affecting pregnancy outcomes. These two applications underscore the versatility of our model across different domains: the first, in Section 6, demonstrates LLDPM’s effectiveness in changepoint detection for time series data, while the second, presented in Section 5.2, illustrates its applicability in a multiview analysis setting.
5.1 Gesture Phase Segmentation data
@@
We present an analysis of video-recorded data for human gesture segmentation. The goal is to segment videos into distinct phases exhibiting different motion patterns, e.g., to identify time lapses of the video that need to be removed from a clip (Parvathy et al., 2021). More specifically, we employ the Gesture Phase Segmentation dataset, publicly accessible for download at the following URL: https://archive.ics.uci.edu/ml/datasets/gesture+phase+segmentation. While Madeo et al. (2013) analyzed this dataset using Support Vector Machines (SVM) to segment, and classify gesture data streams, they did not consider partitioning gesture phases into their components and investigating the temporal dependence between partitions. Our goal, in contrast, is to detect the latent cluster structure at each frame while capturing changepoints, using our proposed LLDPM random partition model with time dependence.
@@
The dataset contains sensor data recordings of users recounting comic book stories facing an Xbox Microsoft KinectTM sensor. The dataset provides scalar velocity and acceleration values over four sensors, placed on the left hand, right hand, left wrist, and right wrist, leading to $n=8$ sensor measurements at regular time intervals (frames). These values were obtained by normalizing hand and wrist positions relative to the head and spine position using a fixed displacement offset of 3 to measure velocity. This offset is a constant adjustment applied to account for the spatial relationship between the sensors and the torso, which helps standardize the measurements and accurately calculate the velocity. Our analysis is based on a processed version of the data, which we prepare following the approach outlined by Hadj-Amar et al. (2024). The preprocessing involves four stages:
@@
Figure 5: Section 5.1, Human Gesture data. Scalar velocity of the left and right hand and the wrists after preprocessing ($T=349$).
@@
Figure 6: Section 5.1, Human Gesture data. Changepoints (dashed line) identified with the LLDPM.
@@
The estimated changepoints are displayed in Figure 6 and appear spread throughout the time series, possibly indicating that the person wearing the sensors is gesticulating while recalling the comic story. When compared with the segmentation in Madeo et al. (2013), the changepoints identified by the LLDPM primarily occur during the activity phase, which correspond to the preparation, retraction, and stroke phases in the video. Additionally, most of these changepoints are observed during transitions between different phases, with a notable increase in frequency during the reading phases. These changepoints likely reflect movements related to the subject’s gesticulation and body language while narrating, as well as shifts in hand and wrist positions during the assigned activity. Figure 7 illustrates the clustering of the eight sensors within a specific time window, chosen for representation purposes. The estimated clustering is obtained by minimizing the lower bound of the posterior expected VI. The numbers on the plot correspond to the sensor measurements. Notably, a distinct pattern allocation emerges, where all the sensors are grouped into a single primary cluster. Additionally, a secondary pattern takes shape, forming two clusters: one containing all the sensors from the left arm (odd numbers) and the other with sensors from the right arm (even numbers) grouped together.
@@
For comparison, we analyzed the same data with four of the alternative models considered in this work. The identified changepoints are displayed in Figures D.14, D.15, D.16 and D.17 in the Supplementary Material. It is notable that partition-based models appear to identify fewer changepoints. For instance, the DRPM detected only 15 changepoints (4.3% of the total time points), primarily at the beginning of the time series, and the LLDPM identified 24 changepoints (6.9%). In contrast, the other alternative models, namely the LDDP, WDDP and GMDDP, identified approximately 14% of the changepoints, with 48, 50 and 51 changepoints, respectively, out of 349 time points. These observations instil confidence in the performance of the LLDPM, which thus might represent a reliable modeling option, particularly in real-world applications where the ground truth is unknown.
@@
Figure 7: Section 5.1, Human Gesture Data. Time-specific clusters obtained by minimizing the lower bound of the posterior expected VI, for the time window from time 335 to time 340. Numbers refer to the measurements of scalar velocity (1-4) and acceleration (5-8) in sensor units placed on the hands and wrists of a subject, whereas colors indicate clusters.
5.2 Collaborative Perinatal Project data
The CPP data were collected to investigate the potential effects of Dichlorodiphenyl-dichloroethylene (DDE) exposure on pregnancy outcomes. In addition to maternal serum DDE concentration, the dataset includes key variables for each enrolled pregnant woman, such as hospital of admission, smoking status, gestational age (in days), and newborn birth weight (in grams). It has been widely used in epidemiological research to examine the impact of prenatal environmental exposures on neonatal health, offering valuable insights for public health policies and maternal care strategies (see, e.g., Klebanoff, 2009). The original study by Longnecker et al. (2001) investigated the relationship between DDE exposure and two primary outcomes: preterm delivery and small-for-gestational-age births.
[…]
To examine variability across hospitals, we perform the same analysis after stratifying the data by hospital. Table 3 reports, for each hospital, the sample size $n$ and the value of $\theta$, which is specified based on the assumption that the prior number of clusters is two for both views. Additionally, it summarizes the results of the LLDPM across hospitals, showing that the estimated probability $\hat{\eta}$ of obtaining different partitions between the two views varies by hospital, ranging from 0.76 to 1. This variability underscores the model’s ability to adapt when applied to subsets of different sizes.
Figure 8: Section 5.2, CPP data. Scatterplot of gestational age (days) and birth weight (grams), with colors indicating cluster assignment.
Hospital 1 2 3 4 5 6 7 8 9 10 11 12 $n$ 481 124 150 77 205 154 141 141 117 384 151 188 $\theta$ 0.32 0.20 0.19 0.22 0.18 0.19 0.19 0.19 0.20 0.16 0.19 0.18 $\hat{\eta}$ 1.00 0.87 1.00 0.77 1.00 0.92 0.99 1.00 1.00 1.00 1.00 1.00
Table 3: Section 5.2, CPP data stratified by hospital. For each of the 12 hospitals in the dataset, we report the sample size $n$, the value of $\theta$, and the estimated probability $\hat{\eta}$ of a change between the latent partitions characterizing the two views: newborn birth weight and gestational age.